-6/((x+1)(x+6))=2+3/(x+6) solve for x , if there is more than 1 solution separate with a comma
Answer by mananth(16946)

-6/((x+1)(x+6))=2+3/(x+6) solve for x , if there is more than 1 solution separate with a comma
-6/((x+1)(x+6))=2+3/(x+6)
-6 / (x+1)(x+6) -3 / (x+6)= 2
-6 - 3(x+1) / (x+1)(x+6) = 2
-6-3x-3 / (x+1)(x+6) =2
-9 -3x = 2( x^2 +7x+6)
-9-3x= 2x^2 +14x+12
2x^2+17x +21=0
2x^2+14x+3x+21=0
2x(x+7) +3(x+7)=0
(x+7)(2x+3)=0
x=-7 , x= -3/2